OpenStudy (anonymous):

Help! please!!!!!!! After a discount of 20%, the price of a pair of shoes, including tax, is $66.91. The following equation can be used to find x, the price of the pair of shoes without the discount: x - 0.2x + 0.06(x - 0.2x) = 66.91 What is the price of the pair of shoes without the discount? $67.76 $71.18 $78.90 $88.04

OpenStudy (anonymous):

@gorv @shinalcantara

OpenStudy (gorv):

add all the term having x

OpenStudy (gorv):

x-0.2x=0.8x

OpenStudy (gorv):

0.8x+0.06*0.8x=66.91

OpenStudy (gorv):

0.8x+0.048x=66.91 0.848x=66.91

OpenStudy (gorv):

x=66.91/0.848

OpenStudy (igreen):

\(x - 0.2x + 0.06(x - 0.2x) = 66.91\) Solve the equation: \(x - 0.2x + 0.06(x - 0.2x) = 66.91\) Distribute: \(x - 0.2x + 0.06x - 0.012x = 66.91\) Subtract: \(0.8x + 0.06x - 0.012x = 66.91\) Add: \(0.86x - 0.012x = 66.91\) Subtract: \(0.848x = 66.91\) Divide 0.848 to both sides: \(x \approx 78.90\)

OpenStudy (anonymous):

thanks!